OpenStudy (anonymous):

Find the area of the indicated region. Enclosed by y = 2^x, y = x + 2, x = −2, and x = 2

OpenStudy (anonymous):

\[\int\limits_{-2}^{-1.69} 2^x-x+2\] and \[\int\limits_{-1.69}^{2} x+2-2^x\] ?

OpenStudy (anonymous):

 what region are you looking for? there should be some inequalities provided

OpenStudy (anonymous):

-1,1

OpenStudy (anonymous):

is it \[y\le2x^2,\quad-2\le x\le1,\quad y\ge x+2\]

OpenStudy (anonymous):

sorry its [-2,2] not -1,1